OpenStudy (anonymous):

A system of equations is shown below: x + y = 3 2x – y = 6 The x-coordinate of the solution to this system of equations is _____. Numerical Answers Expected!

OpenStudy (anonymous):

@am!rah

OpenStudy (anonymous):

@vanorra @lupita1995 @sidsiddhartha @blahhedup101 @HELP!!!!

OpenStudy (anonymous):

@sahrya

OpenStudy (dangerousjesse):

X is equal to three. Actually try working the problem out before you ask.

OpenStudy (anonymous):

Who said I didnt?

OpenStudy (anonymous):

Can you walk me through it to the answer??

OpenStudy (dangerousjesse):

Okay, I'll show you using the process of elimination.

OpenStudy (anonymous):

Alright

OpenStudy (dangerousjesse):

Solve the following system: {x+y = 3 | (equation 1) 2 x-y = 6 | (equation 2) Swap equation 1 with equation 2: {2 x-y = 6 | (equation 1) x+y = 3 | (equation 2) Subtract 1/2 × (equation 1) from equation 2: {2 x-y = 6 | (equation 1) 0 x+(3 y)/2 = 0 | (equation 2) Multiply equation 2 by 2/3: {2 x-y = 6 | (equation 1) 0 x+y = 0 | (equation 2) Add equation 2 to equation 1: {2 x+0 y = 6 | (equation 1) 0 x+y = 0 | (equation 2) Divide equation 1 by 2: {x+0 y = 3 | (equation 1) 0 x+y = 0 | (equation 2) Collect results: \[x = 3\] \[y = 0 \]So do you understand how to do this type of problem now?

OpenStudy (anonymous):

So 3,0?

OpenStudy (anonymous):

And yes, I do understand

OpenStudy (dangerousjesse):

Mhm, and good. Let me know if you need any more help.

OpenStudy (anonymous):

I do need more help on other questions. do you have some time? I have already tried these questions and I just dont get it. can you walk me through? The last question really helped me

OpenStudy (dangerousjesse):

Sure, I have plenty of time.

OpenStudy (anonymous):

The two lines, A and B, are graphed below: Line A is drawn by joining ordered pairs negative 3,18 and 9, negative 6. Line B is drawn by joining ordered pairs negative 5, negative 2 and 8,17 Determine the solution and the reasoning that justifies the solution to the systems of equations. (–4, 6), because both the equations are true for this point (2, 8), because the graph of the two equations intersects at this point (2, 8), because neither of the two equations are true for this coordinate point (–4, 6), because the graph of the two equations intersects the x-axis at these points

OpenStudy (dangerousjesse):